# Hyperparameter Optimization in Machine Learning

### v2

@@ 2.2.1 Estimating generalization
Formalizing this requirement satisfactorily may be a daunting task in certain scenarios (e.g., in generative modelling), but it is easy in the case of supervised learning, where we assume to have access to a task loss :×$\mathcal{L}:\mathcal{H}\times\mathcal{X}\times\mathcal{Y}\to\mathds{R}$. We can define the generalization error of $h_{\lambda}$ as the expected loss:
@@
where $p$ is the unknown task distribution from which the training dataset $\mathcal{D}$ was drawn independently. Widely used metrics such as classification error or mean-squared-error for regression can be interpreted as empirical estimates of (2.6):
@@
where $\mathcal{V}$ is a set of points drawn i.i.d. from $p$, dubbed validation, held-out or development set. For example, the sentiment analysis problem addressed in [yogotama2015-emnlp15] is an instance of supervised binary classification and the task loss is the 00-$1$ error, namely $\mathcal{L}(h_{\lambda},(x,y))=\mathbf{1}\{h_{\lambda}(x)\neq y\}$. Note that in many cases, the internal structure of the learning algorithm, $A(\mathcal{D},\lambda)$, also requires the minimization of a loss function $\mathcal{L}^{\prime}$ with respect to the model’s parameters, which may or may not coincide [^2] For example, in the case of classification, the 0-1 loss is intractable to minimize even for linear classifiers and we usually resort to a convex surrogate [nguyen:2009:surrogate_loss_functions] such as the cross-entropy. with $\mathcal{L}$.
@@
Figure 2.2 shows some examples of other response functions. The leftmost plot concerns a simple ridge regression algorithm to regress median housing prices using the Boston dataset [harrison1978hedonic]. The scalar hyperparameter controls the $L^{2}$ regularization and $\mathcal{L}(h_{\lambda},(x,y))=\|h_{\lambda}(x)-y\|^{2}$. The plot shows the mean loss computed on the training and validation data, respectively. The center plot represent the response function of a two-layer neural network on the same learning scenario. Here the two hyperparameters are an $L^{2}$ regularization coefficient (real-valued, vertical axis) and the number of units in the hidden layer (integer, horizontal axis). The rightmost plot, reproduced from [chen:2018:bayesian_optimization_alphago] represents instead the winning probability of a Go agent introduced in Chapter 1, and somewhat deviates from the standard setting we just discussed.
@@
When estimating generalization of a model, it is essential to compute $\mathscr{M}$ on a set of points that differs from $\mathcal{D}$. In fact, as learning algorithms adapt their hypotheses through $\mathcal{D}$, evaluating $\mathscr{M}(h_{\lambda},\mathcal{D})$ would result in a distorted estimation of (2.6); e.g., underestimation when the model is under-regularized. This is illustrated in Figure 2.2, left. In order to obtain an unbiased estimate of the expected loss (2.6) one needs to draw points from the task distribution $p$ independently both from each other and from $h_{\lambda}$. Otherwise, one carries over unwanted dependencies between training and validation.
@@
In practice, one may split the entirety of the available data into two disjoint sets $\mathcal{D}$ and $\mathcal{V}$ and use the first as input to $A$ and hold out the second for evaluating $f$. 333 Note that a third test set would be needed to estimate generalization capabilities of the hypothesis obtained after HPO! More sophisticated approaches, such as cross-validation, involve splitting the data into $K$ subsets and define the response function as a mean over $K$ evaluation of $\mathscr{M}$, whereby at each round one subset is held out for validation and the rest of the data is used for training [stone1974cross, kohavi:1995:study_cross]. Cross-validation may reduce the variance of the estimate of (2.6) at the expense of increased computational cost.
@@ 2.2.2 Other types of response functions
One may also be interested in optimizing hyperparameters with respect to performance measures not necessarily linked to generalization. For instance, $\mathscr{M}$ could measure the runtime of $A$ when, e.g., tuning optimization hyperparameters [xu2008satzilla, hutter2009paramils], or the inference latency or the memory requirements of $h_{\lambda}$ when looking to deploy models for devices on the edge [benmeziane2021comprehensive]. chen:2018:bayesian_optimization_alphago take $\mathscr{M}$ to be an estimate of the winning rate of the AlphaGo agent that plays with hyperparameters $\lambda$ versus another AlphaGo agent that plays with a fixed baseline hyperparameter configuration $\lambda_{0}$. Figure 2.2 (right) plots the response function (i.e. the winning rate) along two hyperparameter axes.
@@ 2.2.3 General algorithm configuration
Several algorithms, not only those employed in machine learning, require configuration constants that may affect both the quality of the solutions and the computational resources used to find solutions on specific instances [schede:2022:survey_methods_automated]. While in HPO the emphasis is on the former aspect, in general algorithm configuration (AC) the focus is more typically on compute. A typical problem suitable for AC is propositional satisfiability (SAT), where proper choices of parameters can improve solving time significantly [hutter:2009:paramils_automatic_algorithm, hutter_sequential_2011]. Additional interesting examples can be found in the area of discrete optimization, where several parameters may be used to configure metaheuristics [Birattari2009TuningM]. Entering into details would be out of scope for this survey. Still, we would like to point out a striking similarity between AC and HPO. In both cases, optimal (hyper) parameters depend on some unknown probability distribution, such as the data distribution for HPO, or the distribution of future instances of a certain problem in AC. This requires resorting to data (such as a validation set or a set of problem instances) to formulate a proxy empirical optimization problem.
@@ 2.3 Hyperparameter optimization in the literature
Here below we report a few notable milestones and examples of “success stories” of HPO in the literature.
  • bergstra_random_2012, later extended in [bergstra_making_2013], brought the attention of the community to random search as a sensible baseline compared to manual or grid search which, at the time of writing, where widespread approaches to HPO. We cover grid, random and more advanced quasi-random search techniques in Chapter 3.
  • snoek_practical_2012 used Gaussian-process based Bayesian optimization to optimize the hyperparameters of a 3-layer convolutional neural network [krizhevsky-nips12] on CIFAR-10. They improved the manual found hyperparameter by a human expert which marked state-of-the-art at this point by reducing the test error by $3\%$. This papers sparked interest in Bayesian optimization for HPO. Chapter 4 is devoted to model-based HPO and Bayesian optimization in the HPO context.
@@
2.4 Practical desiderata for HPO algorithms
The nested optimization formulation (2.4) abstracts away some aspects which are practically relevant. Together, these characteristics set aside hyperparameter from other optimization scenarios. For this reason one should weigh an HPO algorithm alongside multiple directions. We summarize these desiderata in Table 2.1.
@@ 2.4 Practical desiderata for HPO algorithms
   | Synchronous •  all parallel evaluations need to start at the same time?
   | Asynchronous •  evaluations can be started any time a resource becomes available?
@@
The response function $f(\lambda)$ is typically stochastic, depending on random effects in the computation of $A(\mathcal{D},\lambda)$ (e.g., weight initialization, data ordering, e.g., when performing stochastic gradient descent, etc.). It is also computationally demanding to evaluate. Moreover, HPO is typically only a step in a wider iterative process of model development and refinement. This means that in practice, the emphasis in HPO is to achieve a sufficient decrease of $f(\lambda)$ compared to a reference point (e.g., previously used default configuration) within a given budget $b$ – expressed in time, number of function evaluations, monetary costs, etc. – rather than achieving an (approximate) optimum, which is the typical aim in other branches of optimization. Such a decrease needs to be traded off with compute and runtime costs. A good HPO algorithm should be sample efficient, such that it significantly decreases $f$ with as few evaluations of $f$ as possible. More importantly, it should be wall-clock time efficient, meaning that it finds good solutions as rapidly as possible. While number of evaluations and wall-clock time are related, they are not equivalent. For example, decision-making (i.e., choosing the next configuration to evaluate) counts towards wall-clock time and needs to be balanced against evaluation times of $f$ for model-based HPO algorithms. When comparing different algorithms, time efficiency is captured by plotting the best value of $f$ attained against wall-clock time and computing the area under this curve. Some HPO methods tend to witness rapid initial descent of $f$, but may level off at a sub-optimal value, while others spend more of their initial budget in order to ultimately achieve near-optimal performance.
@@
A powerful principle to increase time efficiency is to consider low-fidelity approximations of the response, which are cheaper to compute. For example, if we train a neural network for 64 epochs, we can access $A_{b}(\mathcal{D},\lambda)$, that is, the model after $b$ epochs of training have elapsed, for free for all $b\leq 64$. Multi-fidelity HPO methods operate on such a sequence of fidelities $f_{b}(\lambda)$, where, in this case, $f(\lambda)=f_{64}(\lambda)$. By controlling the budget $b$ for each trial evaluation $\lambda$ (e.g., stopping training after $b$ epochs), they can often decrease $f(\lambda)$ with a much smaller number of full-fidelity evaluations. These methods will be discussed in detail in Chapter 5. One aspect that is worth noting when choosing an HPO algorithm is if it can leverage low-fidelity evaluation, and how the quality of such evaluation may impact the final result.
@@
We can trade off time efficiency against computation cost by running several evaluations of $f$ in parallel on (say) $W$ workers. For a synchronous system, all $W$ workers start new jobs at the same time. While this reduces wall-clock time, it often leads to idle time when workers need to wait for the slowest job to finish. In an asynchronous system, each worker can be assigned a new job independently from the status of the others. However, HPO decision-making for parallel execution systems comes with additional challenges. In the synchronous case, a batch of $W$ new configurations has to be suggested in one go. Configurations in a batch should be “diverse”, in that they reveal non-redundant information about the optimum [Ginsbourger:10, Desautels:14, Gonzales:16, Wilson2018]. While only a single new configuration needs to be proposed in the asynchronous case, up to $W-1$ responses of running evaluations are not yet available [snoek_practical_2012]. Not taking such pending evaluations into account runs the risk of increased redundancy in the proposed configurations. HPO can also be distributed across several computational nodes that may reside in networked computers. In this case, algorithms should also be parsimonious in terms of communication to reduce overhead as much as possible.
@@
The applicability and accessibility of an HPO algorithm refer to the requirements it may have on the type of problem or the level of access needed to the underlying learning algorithm. Some algorithms may be restricted to certain types of search spaces, e.g., continuous vs integer or categorical domains, or a mix of these. Furthermore, some algorithm can exploit conditionality in the search space, while others cannot. Some HPO methods like gradient-based techniques (Chapter 7) presuppose a white-box access to the underlying learning algorithm in terms of computational structure and/or data processing, while others assume the learning algorithm to be a black box. In between, there are a class of algorithms that require a gray-box access, like the ability to start and stop the execution of the learning algorithm and be able to observe intermediate statistics. Although HPO has been traditionally described and tackled as a black-box optimization problem, more restrictive assumptions may be met in specific cases and HPO algorithm that exploit these assumptions typically show efficiency gains.
Ease of use
The ease of use desideratum refers to the complexity involved in implementing, configuring, and effectively using an HPO algorithm. Some methods may be conceptually simple but require careful tuning of their hyperparameters or auxiliary components. Others may be more complex to implement from scratch or require deep expertise in the underlying techniques (e.g., Bayesian and multi-fidelity optimization) to use them effectively.
2.5 Tackling the search manually
In the next chapters we will introduce the main algorithms for HPO. But in this section we take a few sentences to state the obvious. Searching for the best hyperparameters by hand can be alluring: getting started requires no new frameworks or assumptions, and one may be under the impression that the search could proceed organically, driven by hunches.
Yet manual search can quickly become exhausting. Experienced practitioners may intuitively narrow the search, but novices face a dizzying array of choices. While flexible, tweaking hyperparameters without a systematic approach risks getting lost in an endless maze of configurations. Furthermore, with each trial, insights and analysis of results can guide exploration. While potentially beneficial to speed up the process, this risks overfitting by inadvertently exploiting patterns in the validation data. Manual search also suffers from poor reproducibility, as it is hard to track the steps an user followed to reach some final configuration.
@@ 2.5 Tackling the search manually
We start with grid search, an approach that consists in dividing the search space into a grid of values and querying the response function on each of these values. Next, random search swaps the grid in favour of a fixed, simple distribution over the search space, drawing configurations from such distribution until a budget is exhausted. Finally, we move to quasi-random search strategies, approaches rooted in low-discrepancy sequence generation, in which the simple distributions of random search are replaced by more complex schemes that seek to explore the space more evenly. While grid search is a traditional and classic technique in HPO, consensus has grown in the field that random search should be considered as the baseline method for HPO [bergstra_random_2012]. Quasi-random search may offers benefits over random search, although to date we are not aware of any theoretical result supporting this.

### v3

@@ 2.2.1 Estimating Generalization
Formalizing this requirement satisfactorily may be a daunting task in certain scenarios (e.g., in generative modeling), but it is easy in the case of supervised learning, where we assume to have access to a loss function $\mathcal{L}\mathrel{\mathop{\ordinarycolon}}\mathcal{H}\times\mathcal{X}\times\mathcal{Y}\to\mathds{R}$. We can define the generalization error of $h_{\lambda}$ as the expected loss:
@@
where $p$ is the unknown distribution from which the training dataset $\mathcal{D}$ was drawn independently. Widely used metrics such as classification error or mean-squared-error for regression can be interpreted as empirical estimates of Eq. (2.6):
@@
where $\mathcal{V}$ is a set of points drawn i.i.d. from $p$, dubbed validation, held-out or development set. For example, the sentiment analysis problem addressed by yogotama2015-emnlp15 is an instance of supervised binary classification and the task loss is the 0-11 error, namely $\mathcal{L}(h_{\lambda},(x,y))=\mathds{1}\!\left\{h_{\lambda}(x)\neq y\right\}$. Note that in many cases, the internal structure of the learning algorithm, $A(\mathcal{D},\lambda)$, also requires the minimization of a loss function $\mathcal{J}$ with respect to the model’s parameters, which may or may not coincide with $\mathcal{L}$. For example, in the case of classification, the 0-1 loss is intractable to minimize even for linear classifiers, and we usually resort to a convex surrogate such as the cross-entropy [nguyen:2009:surrogate_loss_functions].
@@
Figure 2.2 shows some examples of other response functions. The leftmost plot concerns a simple ridge regression algorithm to regress median housing prices using the Boston dataset [harrison1978hedonic]. The scalar hyperparameter controls the $L^{2}$ regularization and $\mathcal{L}(h_{\lambda},(x,y))=\|h_{\lambda}(x)-y\|^{2}$. The plot shows the mean loss computed on the training and validation data, respectively. The center plot represents the response function of a two-layer neural network on the same learning scenario. Here the two hyperparameters are an $L^{2}$ regularization coefficient (real-valued, vertical axis) and the number of units in the hidden layer (integer, horizontal axis). The rightmost plot, reproduced from the paper by chen:2018:bayesian_optimization_alphago, represents instead the winning probability of a Go agent introduced in Chapter 1, and somewhat deviates from the standard setting we just discussed.
@@
When estimating generalization of a model, it is essential to compute $\mathscr{M}$ on a set of points that differs from $\mathcal{D}$. In fact, as learning algorithms adapt their hypotheses through $\mathcal{D}$, evaluating $\mathscr{M}(h_{\lambda},\mathcal{D})$ would result in a distorted estimation of Eq. (2.6); e.g., underestimation when the model is under-regularized. This is illustrated in Figure 2.2, left. In order to obtain an unbiased estimate of the expected loss Eq. (2.6) one needs to draw points from the task distribution $p$ independently both from each other and from $h_{\lambda}$. Otherwise, one carries over unwanted dependencies between training and validation.
@@
In practice, one may split the entirety of the available data into two disjoint sets $\mathcal{D}$ and $\mathcal{V}$ and use the first as input to $A$ and hold out the second for evaluating $f$.  [^2] Note that a third test set is needed to estimate generalization capabilities of the hypothesis obtained after HPO! More sophisticated approaches, such as cross-validation, involve splitting the data into $K$ subsets and define the response function as a mean over $K$ evaluation of $\mathscr{M}$, whereby at each round one subset is held out for validation and the rest of the data is used for training [stone1974cross, kohavi:1995:study_cross]. Cross-validation may reduce the variance of the estimate of Eq. (2.6) at the expense of increased computational cost.
@@ 2.2.2 Other Types of Response Functions
One may also be interested in optimizing hyperparameters with respect to performance measures not necessarily linked to generalization. For instance, $\mathscr{M}$ could measure the inference latency or the memory requirements of $h_{\lambda}$ when looking to deploy models for devices on the edge [benmeziane2021comprehensive]. chen:2018:bayesian_optimization_alphago take $\mathscr{M}$ to be an estimate of the winning rate of the AlphaGo agent that plays with hyperparameters $\lambda$ versus another AlphaGo agent that plays with a fixed baseline hyperparameter configuration $\lambda_{0}$. Figure 2.2 (right) plots the response function (i.e. the winning rate) along two hyperparameter axes.
@@ 2.2.3 General Algorithm Configuration
Several algorithms, not only those employed in machine learning, require configuration constants that may affect both the quality of the solutions and the computational resources used to find solutions on specific instances [schede:2022:survey_methods_automated]. While in HPO the emphasis is on the former aspect, in general algorithm configuration (AC) the focus is more typically on compute. A typical problem suitable for AC is propositional satisfiability (SAT), where proper choices of parameters can improve solving time significantly [hutter:2009:paramils_automatic_algorithm, hutter_sequential_2011]. Additional interesting examples can be found in the area of discrete optimization, where several parameters may be used to configure metaheuristics [Birattari2009TuningM]. Entering into details would be out of scope for this survey. Still, we would like to point out a striking similarity between AC and HPO. In both cases, optimal (hyper) parameters depend on some unknown probability distribution, such as the data distribution for HPO, or the distribution of future instances of a certain problem in AC. This requires resorting to data, such as a validation set or a set of problem instances, to formulate a proxy empirical optimization problem.
@@ 2.3 Hyperparameter Optimization in the Literature
Below we report a few notable milestones and examples of “success stories” of HPO in the literature.
  • The paper by bergstra_random_2012, later extended by bergstra_making_2013, brought the attention of the community to random search as a sensible baseline compared to manual or grid search which, at the time of writing, were widespread approaches to HPO. We cover grid, random and more advanced quasi-random search techniques in Chapter 3.
  • snoek_practical_2012 used Gaussian-process based Bayesian optimization to optimize the hyperparameters of a 3-layer convolutional neural network [krizhevsky-nips12] on CIFAR-10. They reduced test error by $3\%$ compared to the previous state-of-the-art that was achieved using manual search. These papers sparked interest in Bayesian optimization for HPO. Chapter 4 is devoted to model-based HPO and Bayesian optimization in the HPO context.
@@
2.4 Practical Considerations for HPO Algorithms
The general formulation of Eq. (2.4) abstracts away specific aspects that are practically important and that collectively set aside HPO from a generic bilevel optimization problem. For this reason, the appraisal of an HPO algorithm should take into account several dimensions, as summarized in Table 2.1.
Table 2.1: List of desirable properties of an HPO algorithm.
@@ 2.4 Practical Considerations for HPO Algorithms
| Synchronous • all parallel evaluations need to start at the same time?
| Asynchronous • evaluations can be started any time a resource becomes available?
@@
The response function $f(\lambda)$ is typically stochastic, depending on random effects in the computation of $A(\mathcal{D},\lambda)$. This may be due for example to weight initialization or to data ordering when performing stochastic gradient descent. It is also computationally demanding to evaluate. Moreover, HPO is typically only a step in a wider iterative process of model development and refinement. This means that in practice, the emphasis in HPO is to achieve a sufficient decrease of $f(\lambda)$ compared to a reference point (e.g., previously used default configuration) within a given budget $b$ — expressed in time, number of function evaluations, monetary costs, etc. — rather than achieving an (approximate) optimum, which is the typical aim in other branches of optimization. Such a decrease needs to be traded off with compute and runtime costs. A good HPO algorithm should be sample efficient, such that it significantly decreases $f$ with as few evaluations of $f$ as possible. More importantly, it should be wall-clock time efficient, meaning that it finds good solutions as rapidly as possible. While number of evaluations and wall-clock time are related, they are not equivalent. For example, decision-making (i.e., choosing the next configuration to evaluate) counts towards wall-clock time and needs to be balanced against evaluation times of $f$ for model-based HPO algorithms. When comparing different algorithms, time efficiency is captured by plotting the best value of $f$ attained against wall-clock time and computing the area under this curve. Some HPO methods tend to witness rapid initial descent of $f$, but may level off at a suboptimal value, while others spend more of their initial budget in order to ultimately achieve near-optimal performance.
@@
A powerful principle to increase time efficiency is to consider low-fidelity approximations of the response, which are cheaper to compute. For example, if we train a neural network for 64 epochs, we can access for free to $b$ checkpoints, for all $b\leq 64$. Multi-fidelity HPO methods operate on such a sequence of fidelities $f_{b}(\lambda)$, where, in this case, $f(\lambda)=f_{64}(\lambda)$. By controlling the budget $b$ for each trial evaluation $\lambda$ (e.g., stopping training after $b$ epochs), they can often decrease $f(\lambda)$ with a much smaller number of full-fidelity evaluations. These methods will be discussed in detail in Chapter 5. One aspect that is worth noting when choosing an HPO algorithm is if it can leverage low-fidelity evaluation, and how the quality of such evaluation may impact the final result.
@@
We can trade off time efficiency against computation cost by running several evaluations of $f$ in parallel on (say) $W$ workers. For a synchronous system, all $W$ workers start new jobs at the same time. While this reduces wall-clock time, it often leads to idle time when workers need to wait for the slowest job to finish. In an asynchronous system, each worker can be assigned a new job independently of the status of the others. However, HPO decision-making for parallel execution systems comes with additional challenges. In the synchronous case, a batch of $W$ new configurations has to be suggested in one go. Configurations in a batch should be “diverse”, in that they reveal non-redundant information about the optimum [Ginsbourger:10, Desautels:14, Gonzales:16, Wilson2018]. While only a single new configuration needs to be proposed in the asynchronous case, up to W1W-1 responses of running evaluations are not yet available [snoek_practical_2012]. Not taking such pending evaluations into account runs the risk of increased redundancy in the proposed configurations. HPO can also be distributed across several computational nodes that may reside in networked computers. In this case, algorithms should also be parsimonious in terms of communication to reduce overhead as much as possible.
@@
The applicability and accessibility of an HPO algorithm refer to the requirements it may have on the type of problem or the level of access needed to the underlying learning algorithm. Some algorithms may be restricted to certain types of search spaces, e.g., continuous vs. integer or categorical domains, or a mix of these. Furthermore, some algorithms can exploit conditionality in the search space, while others cannot. Some HPO methods like gradient-based techniques (Chapter 7) presuppose a white-box access to the underlying learning algorithm in terms of computational structure and/or data processing, while others assume the learning algorithm to be a black box. In between, there are a class of algorithms that require a gray-box access, like the ability to start and stop the execution of the learning algorithm and be able to observe intermediate statistics. Although HPO has been traditionally described and tackled as a black-box optimization problem, more restrictive assumptions may be met in specific cases and HPO algorithm that exploit these assumptions typically show efficiency gains.
Ease of use.
This aspect refers to the complexity involved in implementing, configuring, and effectively using an HPO algorithm. Some methods may be conceptually simple but require careful tuning of their hyperparameters or auxiliary components. Others may be more complex to implement from scratch or require deep expertise in the underlying techniques (e.g., Bayesian and multi-fidelity optimization) to use them effectively.
2.5 Why not Tackling the Search Manually?
In the next chapters we will introduce the main algorithms for HPO. One might question the need for sophisticated algorithms instead of just performing manual search. Indeed, searching for the best hyperparameters by hand can be alluring: getting started requires no new frameworks or assumptions, and one may be under the impression that the search could proceed organically, driven by hunches.
Yet, manual search can quickly become exhausting. Experienced practitioners may intuitively narrow the search, but novices face a dizzying array of choices. While flexible, tweaking hyperparameters without a systematic approach risks getting lost in an endless maze of configurations. Furthermore, with each trial, insights and analysis of results can guide exploration. While potentially beneficial to speed up the process, this risks overfitting by inadvertently exploiting patterns in the validation data. Even for experts, it may be difficult to make valid assumptions about the behavior of a learning algorithm, and it is not always obvious how to carry over expertize from one application domain to another (or from a class of models to another). Finally, manual search also suffers from poor reproducibility, as it is hard to track the steps a user followed to reach some final configuration.
@@ 2.5 Why not Tackling the Search Manually?
We start with grid search, an approach that consists in dividing the search space into a grid of values and querying the response function on each of these values. Next, random search swaps the grid in favor of a fixed, simple distribution over the search space, drawing configurations from such distribution until a budget is exhausted. Finally, we move to quasi-random search strategies, approaches rooted in low-discrepancy sequence generation, in which the simple distributions of random search are replaced by more complex schemes that seek to explore the space more evenly. While grid search is a traditional and classic technique in HPO, consensus has grown in the field that random search should be considered as the baseline method for HPO [bergstra_random_2012]. Quasi-random search may offer benefits over random search, although to date we are not aware of any theoretical result supporting this.